A baloon is at a height of 81m and is ascending upward with a velocity of 12 m/s. A body of 2kg weight is dropped from it. If g =10m/s^2,then body will reach the surface of the earth in how many time?

Grade:9

1 Answers

Vikas TU
7 years ago
h = h0 + vt – 0.5gt^2
Surface will hit when h =0
h0 + vt – 0.5gt^2 = 0
thus,
81 + 12t  – 5t^2 = 0Solutions are t= -2  and t = 5.4 s
thus discarding -3